# Energy Guided Geometric Flow Matching

### v1

@@ 1 Introduction
Generative models are the workhorse of modern AI, enabling us to sample from complex data distributions. Methods based on diffusion have recently excelled at mapping from Gaussian noise distributions to data manifolds, and flow models enable mapping from any noise distribution to the manifold. Typically, these methods are interested in the trajectory that an individual sample takes, from noise to data, only because it enables straightforward simulation with an ODE or SDE solver (yang2024consistency, ). However, for temporal data one may be interested in these trajectories themselves, for example in the context of single-cell trajectory inference.
Recently, more focus has been given to parameterize the trajectories along a manifold, e.g. with fitting expressive Neural ODEs. Early methods focused on the simulation-based setup that required differentiation through an ODE Solver (tong2020trajectorynet, ). However, the improved efficiency of training simulation-free methods such as flow matching (lipman2022flow, ; tong2023improving, ) has granted simpler training to these methods.
@@ 2.1 Denoising Score Matching
The score matching objective (song2020score, ) is the backbone of modern diffusion models, providing an efficient way to learn a family of score functions corresponding to the data distribution under different levels of noise. For a full diffusion model parameterizing density over time, the model takes as input a time variable, but in our setting we consider a monotonic noise schedule of $\sigma_{min}<\dots<\sigma_{max}$ and simply condition on noise. In this case, the score matching loss is given by
@@ 2.2 Conditional Flow Matching
Given two measures, conditional flow matching (tong2023improving, ; lipman2022flow, ) learns a vector field to map one distribution $p_{0}$ to another distribution $p_{1}$. A general form of this loss is given in albergo2023stochastic where one considers a coupling of $\pi$ of the two measures and a parameterization of paths $x_{t}:=x_{t}(x_{0},x_{1})$ that defines a conditional density $p_{t}(\cdot|x_{0},x_{1})$. Then optimizing the loss
@@ 3.1 Flow Matching
Flow matching has rapidly evolved since its original formulations (lipman2022flow, ; albergo2023stochastic, ; liu2022flow, ). In particular, research has focused on incorporating prior knowledge into the the definition of the conditional flows, based on optimal transport (tong2023improving, ; pooladian2023multisample, ), minimal curvature (rohbeck2025modeling, ), and user-defined potential functions (neklyudov2023computational, ), among others. While these methods enrich interpolation flexibility, they rely on Euclidean path parameterization or external priors, rather than directly learning a data-driven metric tensor.
@@ 4.2 Iterative Density Refinement
One shortcoming of score matching is learning the correct normalization of isolated components (wenliang2020blindness, ). Intuitively, this presents a challenge for using the score to induce a metric tensor as even a very accurate score may assign unequal densities to identical but separated components. We consider a novel way of addressing this issue by introducing clustering and stratified sampling (owen2013monte, ).
@@
We use Leiden clustering (traag2019louvain, ) to infer our clusters, as is common in single-cell literature.
@@ 5.2 Synthetic Data
We consider data sampled uniformly on the surface of a sphere in varying dimensions. Note that while this task is handled by RFM (chen2023flow, ) if the manifold is known a priori, learning it directly from the data requires parameterizing the geodesics, and therefore provides a setting to check how accurately our learned geometry matches the ground truth. We omit comparison against CFM, as the optimal solution is simply straight geodesics that ignore the geometry of the sphere entirely.
@@ 5.3.2 Hematopoietic stem cell dataset
We additionally apply our method to a low-dimensional projection of the CITE sequencing dataset (measuring RNA and protein features) of hematopoietic stem cells (open-problems-multimodal, ). This dataset includes four timepoints on days 2, 3, 4, and 7, and therefore we can only evaluate on two heldout timepoints that have a distribution before and after. Our metrics in Table 3 show our method continues to show favorable results.
@@ A.1 Synthetic Experiments
For all our architectures, we use MLPs to parameterize the distance embedding, geodesic and flow networks, and MLPs with residual connections for the score and energy networks. For energy we use the parameterization proposed in[20] and calculate $E(x)=\langle E_{\theta}(x),x\rangle$. We also use skip connections in the embedding network.
@@
We compare against CFM with comparable parameters using the same architectures as well as MFM, which only requires additional parameters for number of clusters $K=2000$ and regularization $\kappa=0.5$, as per[8].
@@ A.2 EB Experiments
For metric flow matching, we used 500 clusters with $\kappa=0.1$. To recreate MFM’s metric tensor, we skip the score and energy network training. We train the embedding and flow networks with identical hyperparameters. For conditional flow matching, the metric tensor is the identity matrix, hence we also skip the score and energy networks. The rest of the model is trained identically. The data is available at https://data.mendeley.com/datasets/hhny5ff7yj/1 and preprocessed according to the same code as[22].

### v2

@@ 1 Introduction
Generative models are the workhorse of modern AI, enabling us to sample from complex data distributions. Methods based on diffusion have recently excelled at mapping from Gaussian noise distributions to data manifolds, and flow models enable mapping from any noise distribution to the manifold. Typically, these methods are interested in the trajectory that an individual sample takes, from noise to data, only because it enables straightforward simulation with an ODE or SDE solver (yang2024consistency). However, for temporal data one may be interested in these trajectories themselves, for example in the context of single-cell trajectory inference.
Recently, more focus has been given to parameterize the trajectories along a manifold, e.g. with fitting expressive Neural ODEs. Early methods focused on the simulation-based setup that required differentiation through an ODE Solver (tong2020trajectorynet). However, the improved efficiency of training simulation-free methods such as flow matching (lipman2022flow; tong2023improving) has granted simpler training to these methods.
@@ 2.1 Denoising Score Matching
The score matching objective (song2020score) is the backbone of modern diffusion models, providing an efficient way to learn a family of score functions corresponding to the data distribution under different levels of noise. For a full diffusion model parameterizing density over time, the model takes as input a time variable, but in our setting we consider a monotonic noise schedule of $\sigma_{min}<\dots<\sigma_{max}$ and simply condition on noise. In this case, the score matching loss is given by
@@ 2.2 Conditional Flow Matching
Given two measures, conditional flow matching (tong2023improving; lipman2022flow) learns a vector field to map one distribution $p_{0}$ to another distribution $p_{1}$. A general form of this loss is given in albergo2023stochastic where one considers a coupling of $\pi$ of the two measures and a parameterization of paths $x_{t}:=x_{t}(x_{0},x_{1})$ that defines a conditional density $p_{t}(\cdot|x_{0},x_{1})$. Then optimizing the loss
@@ 3.1 Flow Matching
Flow matching has rapidly evolved since its original formulations (lipman2022flow; albergo2023stochastic; liu2022flow). In particular, research has focused on incorporating prior knowledge into the the definition of the conditional flows, based on optimal transport (tong2023improving; pooladian2023multisample), minimal curvature (rohbeck2025modeling), and user-defined potential functions (neklyudov2023computational), among others. While these methods enrich interpolation flexibility, they rely on Euclidean path parameterization or external priors, rather than directly learning a data-driven metric tensor.
@@ 4.2 Iterative Density Refinement
One shortcoming of score matching is learning the correct normalization of isolated components (wenliang2020blindness). Intuitively, this presents a challenge for using the score to induce a metric tensor as even a very accurate score may assign unequal densities to identical but separated components. We consider a novel way of addressing this issue by introducing clustering and stratified sampling (owen2013monte).
@@
We use Leiden clustering (traag2019louvain) to infer our clusters, as is common in single-cell literature.
@@ 5.2 Synthetic Data
We consider data sampled uniformly on the surface of a sphere in varying dimensions. Note that while this task is handled by RFM (chen2023flow) if the manifold is known a priori, learning it directly from the data requires parameterizing the geodesics, and therefore provides a setting to check how accurately our learned geometry matches the ground truth. We omit comparison against CFM, as the optimal solution is simply straight geodesics that ignore the geometry of the sphere entirely.
@@ 5.3.2 Hematopoietic stem cell dataset
We additionally apply our method to a low-dimensional projection of the CITE sequencing dataset (measuring RNA and protein features) of hematopoietic stem cells (open-problems-multimodal). This dataset includes four timepoints on days 2, 3, 4, and 7, and therefore we can only evaluate on two heldout timepoints that have a distribution before and after. Our metrics in Table 3 show our method continues to show favorable results.
@@ 6 Conclusion
Acknowledgments
This work was made possible by support from the MacMillan Family and the MacMillan Center for the Study of the Non-Coding Cancer Genome at the New York Genome Center. A.Z. is the Sijbrandij Foundation Quantitative Biology Fellow of the Damon Runyon Cancer Research Foundation (DRQ-26-25).
@@ A.1 Synthetic Experiments
For all our architectures, we use MLPs to parameterize the distance embedding, geodesic and flow networks, and MLPs with residual connections for the score and energy networks. For energy we use the parameterization proposed in[thornton2025composition] and calculate $E(x)=\langle E_{\theta}(x),x\rangle$. We also use skip connections in the embedding network.
@@
We compare against CFM with comparable parameters using the same architectures as well as MFM, which only requires additional parameters for number of clusters $K=2000$ and regularization $\kappa=0.5$, as per[kapusniak2024metric].
@@ A.2 EB Experiments
For metric flow matching, we used 500 clusters with $\kappa=0.1$. To recreate MFM’s metric tensor, we skip the score and energy network training. We train the embedding and flow networks with identical hyperparameters. For conditional flow matching, the metric tensor is the identity matrix, hence we also skip the score and energy networks. The rest of the model is trained identically. The data is available at https://data.mendeley.com/datasets/hhny5ff7yj/1 and preprocessed according to the same code as[tong2020trajectorynet].